OpenStudy (anonymous):

write a polynomial function in standard form with zero's at -1,-4 and -3 can you please explain the steps?

OpenStudy (anonymous):

(x+1)(x+4)(x+3) is I believe what you want

OpenStudy (anonymous):

this expands to x^3 +8x^2 +19x + 12

OpenStudy (anonymous):

yes thats right but please show me how to expand it

OpenStudy (anonymous):

The graph of this function passes through the x-axis at the points where the polynomial = 0 ie whwn x = -1, -3 and -4. OK first deal with (x+1)(x+4): first multiply (x+4 by x giving x^2 + 4x then multiply (x + 4) by +1 giving x + 4 Adding these together gives x^2 + 4x + x +4 = x^2 + 5x + 4 Finally we multiply x^2 + 5x + 4 by the other term (x + 3): as before first multiply by x then by 3 then add the results together does this explain it ok? If not get back to me

OpenStudy (anonymous):

does it matter what we deal with first? why (x+1)(x+4) and not (X+1)(x+3)

OpenStudy (anonymous):

no you'll still get the same result